Part A) You have been managing a $5 million portfolio that has a beta of 1.25 and a required rate of return of 8.875%. The current risk-free rate is 2%. Assume that you receive another $500,000. If you invest the money in a stock with a beta of 1.45, what will be the required return on your $5.5 million portfolio? Do not round intermediate calculations. Round your answer to two decimal places.

Part B) Carnes Cosmetics Co.'s stock price is $54, and it recently paid a $2.50 dividend. This dividend is expected to grow by 25% for the next 3 years, then grow forever at a constant rate, g; and rs = 15%. At what constant rate is the stock expected to grow after Year 3? Do not round intermediate calculations. Round your answer to two decimal places.
Answer #1

1.

Calculating Market Rate by using CAPM model,

0.08875 = 0.02 + 1.25(Rm - 0.02)

Rm = 7.50%

New Beta = [5,000,000(1.25) + 500,000(1.45)]/5500,000

New Beta = 1.268

Required Return = 0.02 + 1.268(0.075 - 0.02)

Required Return = 8.97%

2.

54 = 2.50(1.25/1.15) + 2.50(1.25/1.15)2 + 2.50(1.25/1.15)3 + X/(1.15)3

54 = 2.72 + 2.95 + 3.21 + X/(1.15)3

X = $68.79

So,

68.79 = 2.50(1.25)3(1 + g)/(0.15 - g)

14.10(0.15 - g) = (1 + g)

2.115 - 14.10g = 1 + g

g = 7.38%